# Fodor space in generalized descriptive set theory

### 10 Mar 2026 v1

@@ 2. Fodor space
In the $n$-th move, first $\bf I$} chooses an ordinal $\beta_{n}<\alpha$ such that $X_{\beta_{n}}\subset\alpha$ and $X_{\beta_{n-1}}\subseteq X_{\beta_{n}}$. Then $\bf II$} chooses an ordinal $\theta_{n}<\alpha$ such that $dom(f_{\theta_{n}}),ran(f_{\theta_{n}})\subset\alpha$, $X_{\beta_{n}}\subseteq dom(f_{\theta_{n}})\cap ran(f_{\theta_{n}})$ and $f_{\theta_{n-1}}\subseteq f_{\theta_{n}}$ (if $n=0$ then $X_{\beta_{n-1}}=\emptyset$ and $f_{\theta_{n-1}}=\emptyset$).
@@ 3. Trees
We will construct a general version of Hyttinen-Kulikov’s coloured trees, this will allow us to construct models from $\eta\in\mathbb{F}$.
@@
For all $f\in 2^{\kappa}$ and stationary $E\subseteq\operatorname{cf}(\lambda)$, define the tree $(\mathcal{R}^{E}_{f},r^{E}_{f})$ as, $\mathcal{R}^{E}_{f}$ the set of all strictly increasing functions, $\eta$, from some $\alpha\leq\lambda\cdot\lambda$ to $\kappa$, $\lambda$-cofinal to $E$, and if $\eta$ has domain $\lambda\cdot\lambda$, then $r^{E}_{f}(\eta)=f(sup(rng(\eta)))$.
@@
      there is no such ordinal $m$ and then $c^{E}_{H}(\eta):=H(sup(rang(\eta_{5})))$.
@@
Since $\eta_{5}$ is increasing and $sup(\operatorname{rng}(\eta_{3}))\geq sup(\operatorname{rng}(\eta_{5}))\geq sup(\operatorname{rng}(\eta_{2}))$, $sup(\operatorname{rng}(\eta_{2}))\geq sup(\operatorname{rng}(\eta_{3}))$ and $sup(\operatorname{rng}(\eta_{2}))\geq sup(\operatorname{rng}(\eta_{4}))$, this leads us to
$sup(\operatorname{rng}(\eta_{4}))\leq sup(\operatorname{rng}(\eta_{3}))=sup(\operatorname{rng}(\eta_{5}))=sup(\operatorname{rng}(\eta_{2})).$
@@
Suppose $\xi\in(I_{\sigma}^{f})^{\alpha}$ and $\eta\in I_{\sigma}^{f}$. If $dom(\xi)$ is a successor ordinal smaller than $\lambda\cdot\lambda$, $\xi\subsetneq\eta$ and for every $k$ in $dom(\eta)\backslash dom(\xi)$, $\eta_{1}(k)=\xi_{1}(max(dom(\xi)))$ and $\eta_{1}(k)>0$, then $\eta\in(I_{\sigma}^{f})^{\alpha}$.
@@
For all $f:\kappa\rightarrow\kappa$ and $\sigma<\kappa$, we will use $I^{f}_{\sigma}$ to construct $A^{f}_{\sigma}$, an element of $K_{tr}^{\lambda\cdot\lambda}$. Let us denote by $\operatorname{acc}(\kappa)=\{\alpha<\kappa\mathrel{|}\allowbreak\alpha=0~\text{or}~\alpha\text{ is a limit ordinal}\}$. For all $\alpha\in\operatorname{acc}(\kappa)$ and $\eta\in(I_{\sigma}^{f})^{\alpha}$ with $dom(\eta)=m<\lambda\cdot\lambda$ define
$W_{\eta}^{\alpha}=\{\zeta\mathrel{|}\allowbreak dom(\zeta)=[m,s),m\leq s\leq\lambda\cdot\lambda,\eta^{\frown}\zeta\in(I_{\sigma}^{f})^{\alpha+\omega},\eta^{\frown}\langle m,\zeta(m)\rangle\notin(I_{\sigma}^{f})^{\alpha}\}.$
@@
Let us denote by $\mathbb{T}=(\kappa\times\omega\times\operatorname{acc}(\kappa)\times(\lambda\cdot\lambda)\times\kappa\times\kappa\times\kappa\times\kappa)^{\leq\lambda\cdot\lambda}$. For every $\xi\in\mathbb{T}$ there are functions $\{\xi_{i}\in\kappa^{\leq\lambda\cdot\lambda}\mathrel{|}\allowbreak 0<i\leq 8\}$ such that for all $i\leq 8$, $dom(\xi_{i})=dom(\xi)$ and for all $n\in dom(\xi)$, $\xi(n)=(\xi_{1}(n),\xi_{2}(n),\xi_{3}(n),\xi_{4}(n),\xi_{5}(n),\xi_{6}(n),\xi_{7}(n),\xi_{8}(n))$. For every $\xi\in\mathbb{T}$ let us denote $(\xi_{4},\xi_{5},\xi_{6},\xi_{7},\xi_{8})$ by $\overline{\xi}$.
@@
For all $\alpha\in\operatorname{acc}(\kappa)$ and $\eta\in\mathbb{T}$ with $\overline{\eta}\in I_{\sigma}^{f}$ , $dom(\eta)=m<\lambda\cdot\lambda$ define $\Gamma_{\eta}^{\alpha}$ as follows:
@@
    $\overline{\xi}\mathbin{\upharpoonright}dom(\xi)\backslash m\in W^{\alpha}_{\eta},$
@@
    for all $n\in dom(\xi)\backslash m$, let $\xi_{2}(n)$ be the unique $r<\omega$ such that $\varsigma_{\zeta}^{\alpha}(r)=\overline{\xi}(n)$, where $\zeta=\overline{\xi}\mathbin{\upharpoonright}n$.
@@
For $\eta\in\mathbb{T}$ with $\overline{\eta}\in I_{\sigma}^{f}$, $dom(\eta)=m<\lambda\cdot\lambda$ define
@@
$\mathbb{I}_{\sigma}^{f}(n+1)=\mathbb{I}_{\sigma}^{f}(n)\cup\bigcup_{\eta\in\mathbb{I}_{\sigma}^{f}(n)~dom(\eta)=n}\Gamma(\eta).$
@@
$\mathbb{I}_{\sigma}^{f}(n)=\bar{\mathbb{I}}_{\sigma}^{f}(n)\cup\{\eta\in\mathbb{T}\mathrel{|}\allowbreak dom(\eta)=n\ \&\ \forall m<n\ (\eta\mathbin{\upharpoonright}m\in\bar{\mathbb{I}}_{\sigma}^{f}(n))\ \&\ \overline{\eta}\in I_{\sigma}^{f}\}.$
@@
$(A^{f}_{\sigma})^{\alpha}=\{\eta\in A^{f}_{\sigma}\mathrel{|}\allowbreak rng(\eta)\subseteq\vartheta\times\omega\times\vartheta\times(\lambda\cdot\lambda)\times\vartheta^{4}\text{ for some }\vartheta<\alpha\}.$
@@
$T_{\sigma}^{f}=A^{f}_{\sigma}\backslash\{\eta\in A^{f}_{\sigma}\mathrel{|}\allowbreak dom(\eta)=\lambda\cdot\lambda\ \&\ d^{f}_{\sigma}(\eta)=0\}$
@@
Clearly (2) follows from the way we chose $\tau$.
@@ 4. Ehrenfeucht-Mostowski models
In the case of stable unsuperstable theories, the construction of the models should follow the generalized Ehrenfeucht-Mostowski models (see [11] Theorem 1.3 and [12] Chapter VII). The construction of such models depends on the existence of a proper function $\Phi$. Unfortunately, the argument in [12] for the existence of a proper function, fails when the ordered tree has high $\omega\cdot\omega$ or higher.
@@
    $\nu_{i}\in(T^{f})^{\bar{\delta}}$ for some $\bar{\delta}>\delta$ and there is $\bar{\tau}\leq dom(\nu_{i})$ such that $\nu_{i}\mathbin{\upharpoonright}\bar{\tau}\in\operatorname{acc}((T^{f})^{\delta})\setminus(T^{f})^{\delta}$. There is $k\leq 8$ $\sup(\operatorname{rng}((\nu_{i}\mathbin{\upharpoonright}\bar{\tau})_{k}))=\delta$. By Remark 3.4, $\sup(\operatorname{rng}((\nu_{i}\mathbin{\upharpoonright}\bar{\tau})_{8}))=\delta$. Let $B_{i}:=\operatorname{acc}(\operatorname{rng}((\nu_{i})_{8}\mathbin{\upharpoonright}\bar{\tau}))$.
@@
For all $i<n$ satisfying Clause (2) and $\bar{\tau}\leq dom(\nu_{i})$ witnessing it, $\sup(\operatorname{rng}((\nu_{i})_{8}\mathbin{\upharpoonright}\bar{\tau}))=\delta$. Since $\delta\in S^{g}$, $\delta\notin S_{1}$ and $\bar{\tau}=dom(\nu_{i})=\lambda\cdot\lambda$. Thus $\nu_{i}\in\mathcal{B}(T^{f}_{\sigma_{f}})$. Recall that $\sigma_{f}\neq\sigma_{g}$. Thus, by the construction of $T^{f}$ specifically Definition 3.3 clause (8),
@@
It is clear that $\langle B^{f}(\eta,\alpha)\mathrel{|}\allowbreak\alpha<\kappa\rangle$ is a filtration of $Suc_{T^{f}}(\eta)$. By Theorem 3.17, $T^{f}$ is $(\kappa,\varepsilon)$-nice, in particular $Suc_{T^{f}}(\eta)$ is isomorphic to $\mathcal{I}$. Since any two representations coincide in a club, for any $\eta\in T^{f}$ there is a club $C_{\eta}$ such that for all $\delta\in C_{\eta}$ with $cf(\delta)\geq\varepsilon$ and $\nu\in Suc_{T^{f}}(\eta)$ there is $\beta<\delta$ such that
@@
$\bar{C}^{f}=\{\delta<\kappa\mathrel{|}\allowbreak cf(\delta)\geq\varepsilon\text{ and for all }\eta\in(T^{f})_{\delta},\ \delta\in C_{\eta}\}$
and $C^{f}$ be $\bar{C}^{f}$ closed under $\alpha$-limits for $\alpha<\varepsilon$. Notice that $C^{f}$ is a club that satisfies that for all $\delta\in C^{f}$ with $cf(\delta)\geq\varepsilon$, $\eta\in T^{f}$, $T^{f}\not\models P_{\lambda\cdot\lambda}(\eta)$, and $\nu\in Suc_{T^{f}}(\eta)$ there is $\beta<\delta$ such that,
@@
  • $C_{2}=\{\delta\in C^{\prime}_{2}\mathrel{|}\allowbreak cf(\delta)\geq\lambda\}$
@@ 5. Conclusions
For every $f\in\mathbb{F}^{*}$ there is a bijection $E_{f}:dom(\mathcal{M}_{f})\rightarrow\kappa$, such that for every $f,g\in\mathbb{F}^{*}$ and $\alpha<\kappa$, if $f\mathbin{\upharpoonright}\alpha=g\mathbin{\upharpoonright}\alpha$, then $E_{f}\mathbin{\upharpoonright}dom(\mathcal{M}_{f,\alpha})=E_{g}\mathbin{\upharpoonright}dom(\mathcal{M}_{g,\alpha})$ (see [7]). Let us denote by $\pi$ the bijection from Definition 1.4, define the function $\mathcal{G}$ by:
@@
To show that $G:\mathbb{F}\rightarrow 2^{\kappa}$, $G(f)=\mathcal{G}(\mathcal{M}_{f})$ is continuous, let $\zeta\mathbin{\upharpoonright}\alpha]$ be a basic open set and $\xi\in G^{-1}[[\zeta\mathbin{\upharpoonright}\alpha]]$. There is $\beta<\kappa$ such that for all $\epsilon<\alpha$, if $\epsilon=\pi(m,a_{1},\ldots,a_{n})$, then $E^{-1}_{\xi}(a_{i})\in dom(\mathcal{M}_{\xi,\beta})$ holds for all $i\leq n$. Since for all $\eta\in[\xi\mathbin{\upharpoonright}\beta]$ it holds that $\mathcal{M}_{\eta,\beta}=\mathcal{M}_{\xi,\beta}$, for any $\epsilon<\alpha$ that satisfies $\epsilon=\pi(m,a_{1},\ldots,a_{n})$

### 23 May 2026 v2

@@ 1. Introduction
Notice that strictly stable (stable unsuperstable) theories are not covered in the previous theorem. As it was showed in [4] OCP implies that the theory is unsuperstable, thus stable theories with OCP are strictly stable. Unfortunately not all strictly stable theories have the OCP. In a private correspondence with the second author; Mennuni pointed out that in a non-multi-dimensional theory, no type is orthogonal to any set (see Lemma V 5.4 [15]). Thus a strictly stable non-multi-dimensional theory doesn’t have the OCP. On the other hand any theory of modules is non-multi-dimensional (see Corollary 3.5 [12] or Corollary 6.21 [13]). So any strictly stable theory of modules (e.g. the theory of $\mathbb{Z}^{\aleph_{0}}$ with the Abelian group structure) doesn’t have the OCP (this answers Question 4.5 [7], originally asked by Villaveces).
@@ 2. Fodor space
In the $n$-th move, first $\bf I$} chooses an ordinal $\beta_{n}<\alpha$ such that $X_{\beta_{n}}\subset\alpha$ and $X_{\beta_{n-1}}\subseteq X_{\beta_{n}}$. Then $\bf II$} chooses an ordinal $\theta_{n}<\alpha$ such that $\operatorname{dom}(f_{\theta_{n}}),\operatorname{rng}(f_{\theta_{n}})\subset\alpha$, $X_{\beta_{n}}\subseteq\operatorname{dom}(f_{\theta_{n}})\cap\operatorname{rng}(f_{\theta_{n}})$ and $f_{\theta_{n-1}}\subseteq f_{\theta_{n}}$ (if $n=0$ then $X_{\beta_{n-1}}=\emptyset$ and $f_{\theta_{n-1}}=\emptyset$).
@@ 3. Trees
We will construct a variation of Hyttinen-Kulikov’s coloured trees, this will allow us to construct models from $\eta\in\mathbb{F}$.
@@
For all $f\in 2^{\kappa}$ and stationary $E\subseteq\operatorname{cf}(\lambda)$, define the tree $(\mathcal{R}^{E}_{f},r^{E}_{f})$ as, $\mathcal{R}^{E}_{f}$ the set of all strictly increasing functions, $\eta$, from some $\alpha\leq\lambda\cdot\lambda$ to $\kappa$, $\lambda$-cofinal to $E$, and if $\eta$ has domain $\lambda\cdot\lambda$, then $r^{E}_{f}(\eta)=f(\sup(\operatorname{rng}(\eta)))$.
@@
      there is no such ordinal $m$ and then $c^{E}_{H}(\eta):=H(\sup(\operatorname{rng}(\eta_{5})))$.
@@
Since $\eta_{5}$ is increasing and $\sup(\operatorname{rng}(\eta_{3}))\geq\sup(\operatorname{rng}(\eta_{5}))\geq\sup(\operatorname{rng}(\eta_{2}))$, $\sup(\operatorname{rng}(\eta_{2}))\geq\sup(\operatorname{rng}(\eta_{3}))$ and $\sup(\operatorname{rng}(\eta_{2}))\geq\sup(\operatorname{rng}(\eta_{4}))$, this leads us to
$\sup(\operatorname{rng}(\eta_{4}))\leq\sup(\operatorname{rng}(\eta_{3}))=\sup(\operatorname{rng}(\eta_{5}))=\sup(\operatorname{rng}(\eta_{2})).$
@@
[…]
@@
For all $f:\kappa\rightarrow\kappa$ and $\sigma<\kappa$, we will use $I^{f}_{\sigma}$ to construct $A^{f}_{\sigma}$, an element of $K_{tr}^{\lambda\cdot\lambda}$. Let us denote by $\operatorname{acc}(\kappa)=\{\alpha<\kappa\mathrel{|}\allowbreak\alpha=0~\text{or}~\alpha\text{ is a limit ordinal}\}$. For all $\alpha\in\operatorname{acc}(\kappa)$ and $\eta\in(I_{\sigma}^{f})^{\alpha}$ with $\operatorname{dom}(\eta)=m<\lambda\cdot\lambda$ define
$W_{\eta}^{\alpha}=\{\zeta\mathrel{|}\allowbreak\operatorname{dom}(\zeta)=[m,s),m\leq s\leq\lambda\cdot\lambda,\eta^{\frown}\zeta\in(I_{\sigma}^{f})^{\alpha+\omega},\eta^{\frown}\langle m,\zeta(m)\rangle\notin(I_{\sigma}^{f})^{\alpha}\}.$
@@
Let us denote by $\mathbb{T}=(\kappa\times\omega\times\operatorname{acc}(\kappa)\times(\lambda\cdot\lambda)\times\kappa\times\kappa\times\kappa\times\kappa)^{\leq\lambda\cdot\lambda}$. For every $\xi\in\mathbb{T}$ there are functions $\{\xi_{i}\in\kappa^{\leq\lambda\cdot\lambda}\mathrel{|}\allowbreak 0<i\leq 8\}$ such that for all $i\leq 8$, $\operatorname{dom}(\xi_{i})=\operatorname{dom}(\xi)$ and for all $n\in\operatorname{dom}(\xi)$, $\xi(n)=(\xi_{1}(n),\xi_{2}(n),\xi_{3}(n),\xi_{4}(n),\xi_{5}(n),\xi_{6}(n),\xi_{7}(n),\xi_{8}(n))$. For every $\xi\in\mathbb{T}$ let us denote $(\xi_{4},\xi_{5},\xi_{6},\xi_{7},\xi_{8})$ by $\overline{\xi}$.
@@
For all $\alpha\in\operatorname{acc}(\kappa)$ and $\eta\in\mathbb{T}$ with $\overline{\eta}\in I_{\sigma}^{f}$ , $\operatorname{dom}(\eta)=m<\lambda\cdot\lambda$ define $\Gamma_{\eta}^{\alpha}$ as follows:
@@
    $\overline{\xi}\mathbin{\upharpoonright}\operatorname{dom}(\xi)\backslash m\in W^{\alpha}_{\eta},$
@@
[…]
@@
For $\eta\in\mathbb{T}$ with $\overline{\eta}\in I_{\sigma}^{f}$, $\operatorname{dom}(\eta)=m<\lambda\cdot\lambda$ define
@@
$\mathbb{I}_{\sigma}^{f}(n+1)=\mathbb{I}_{\sigma}^{f}(n)\cup\bigcup_{\eta\in\mathbb{I}_{\sigma}^{f}(n)~\operatorname{dom}(\eta)=n}\Gamma(\eta).$
@@
$\mathbb{I}_{\sigma}^{f}(n)=\bar{\mathbb{I}}_{\sigma}^{f}(n)\cup\{\eta\in\mathbb{T}\mathrel{|}\allowbreak\operatorname{dom}(\eta)=n\ \&\ \forall m<n\ (\eta\mathbin{\upharpoonright}m\in\bar{\mathbb{I}}_{\sigma}^{f}(n))\ \&\ \overline{\eta}\in I_{\sigma}^{f}\}.$
@@
$(A^{f}_{\sigma})^{\alpha}=\{\eta\in A^{f}_{\sigma}\mathrel{|}\allowbreak\operatorname{rng}(\eta)\subseteq\vartheta\times\omega\times\vartheta\times(\lambda\cdot\lambda)\times\vartheta^{4}\text{ for some }\vartheta<\alpha\}.$
@@
$T_{\sigma}^{f}=A^{f}_{\sigma}\backslash\{\eta\in A^{f}_{\sigma}\mathrel{|}\allowbreak\operatorname{dom}(\eta)=\lambda\cdot\lambda\ \&\ d^{f}_{\sigma}(\eta)=0\}$
@@
Clearly (2) follows from the way we chose $\tau$. ∎
@@ 4. Ehrenfeucht-Mostowski models
The construction of Ehrenfeucht-Mostowski models for stable unsuperstable theories is different than for the other cases. The construction of the models follow the generalized Ehrenfeucht-Mostowski models (see [14] Theorem 1.3 and [15] Chapter VII). The construction of such models depends on the existence of a proper function $\Phi$. Unfortunately, the argument in [15] for the existence of a proper function, fails when the ordered tree has high $\omega\cdot\omega$ or higher.
@@
    $\nu_{i}\in(T^{f})^{\bar{\delta}}$ for some $\bar{\delta}>\delta$ and there is $\bar{\tau}\leq\operatorname{dom}(\nu_{i})$ such that $\nu_{i}\mathbin{\upharpoonright}\bar{\tau}\in\operatorname{acc}((T^{f})^{\delta})\setminus(T^{f})^{\delta}$. There is $k\leq 8$ $\sup(\operatorname{rng}((\nu_{i}\mathbin{\upharpoonright}\bar{\tau})_{k}))=\delta$. By Remark 3.4, $\sup(\operatorname{rng}((\nu_{i}\mathbin{\upharpoonright}\bar{\tau})_{8}))=\delta$. Let $B_{i}:=\operatorname{acc}(\operatorname{rng}((\nu_{i})_{8}\mathbin{\upharpoonright}\bar{\tau}))$.
@@
For all $i<n$ satisfying Clause (2) and $\bar{\tau}\leq\operatorname{dom}(\nu_{i})$ witnessing it, $\sup(\operatorname{rng}((\nu_{i})_{8}\mathbin{\upharpoonright}\bar{\tau}))=\delta$. Since $\delta\in S^{g}$, $\delta\notin S_{1}$ and $\bar{\tau}=\operatorname{dom}(\nu_{i})=\lambda\cdot\lambda$. Thus $\nu_{i}\in\mathcal{B}(T^{f}_{\sigma_{f}})$. Recall that $\sigma_{f}\neq\sigma_{g}$. Thus, by the construction of $T^{f}$ specifically Definition 3.3 clause (8),
@@
It is clear that $\langle B^{f}(\eta,\alpha)\mathrel{|}\allowbreak\alpha<\kappa\rangle$ is a filtration of $Suc_{T^{f}}(\eta)$. By Theorem 3.17, $T^{f}$ is $(\kappa,\varepsilon)$-nice, in particular $Suc_{T^{f}}(\eta)$ is isomorphic to $\mathcal{I}$. Since any two representations coincide in a club, for any $\eta\in T^{f}$ there is a club $C_{\eta}$ such that for all $\delta\in C_{\eta}$ with $\operatorname{cf}(\delta)\geq\varepsilon$ and $\nu\in Suc_{T^{f}}(\eta)$ there is $\beta<\delta$ such that
@@
$\bar{C}^{f}=\{\delta<\kappa\mathrel{|}\allowbreak\operatorname{cf}(\delta)\geq\varepsilon\text{ and for all }\eta\in(T^{f})_{\delta},\ \delta\in C_{\eta}\}$
and $C^{f}$ be $\bar{C}^{f}$ closed under $\alpha$-limits for $\alpha<\varepsilon$. Notice that $C^{f}$ is a club that satisfies that for all $\delta\in C^{f}$ with $\operatorname{cf}(\delta)\geq\varepsilon$, $\eta\in T^{f}$, $T^{f}\not\models P_{\lambda\cdot\lambda}(\eta)$, and $\nu\in Suc_{T^{f}}(\eta)$ there is $\beta<\delta$ such that,
@@
  • $C_{2}=\{\delta\in C^{\prime}_{2}\mathrel{|}\allowbreak\operatorname{cf}(\delta)\geq\lambda\}$
@@ 5. Conclusions
For every $f\in\mathbb{F}^{*}$ there is a bijection $E_{f}:\operatorname{dom}(\mathcal{M}_{f})\rightarrow\kappa$, such that for every $f,g\in\mathbb{F}^{*}$ and $\alpha<\kappa$, if $f\mathbin{\upharpoonright}\alpha=g\mathbin{\upharpoonright}\alpha$, then $E_{f}\mathbin{\upharpoonright}\operatorname{dom}(\mathcal{M}_{f,\alpha})=E_{g}\mathbin{\upharpoonright}\operatorname{dom}(\mathcal{M}_{g,\alpha})$ (see [8]). Let us denote by $\pi$ the bijection from Definition 1.4, define the function $\mathcal{G}$ by:
@@
To show that $G:\mathbb{F}\rightarrow 2^{\kappa}$, $G(f)=\mathcal{G}(\mathcal{M}_{f})$ is continuous, let $\zeta\mathbin{\upharpoonright}\alpha]$ be a basic open set and $\xi\in G^{-1}[[\zeta\mathbin{\upharpoonright}\alpha]]$. There is $\beta<\kappa$ such that for all $\epsilon<\alpha$, if $\epsilon=\pi(m,a_{1},\ldots,a_{n})$, then $E^{-1}_{\xi}(a_{i})\in\operatorname{dom}(\mathcal{M}_{\xi,\beta})$ holds for all $i\leq n$. Since for all $\eta\in[\xi\mathbin{\upharpoonright}\beta]$ it holds that $\mathcal{M}_{\eta,\beta}=\mathcal{M}_{\xi,\beta}$, for any $\epsilon<\alpha$ that satisfies $\epsilon=\pi(m,a_{1},\ldots,a_{n})$